### v1

@@
Abstract
Despite being trained on extensive datasets, current computer vision systems lag behind human children in learning about the visual world. One possible reason for this discrepancy is the fact that humans actively explore their environment as embodied agents, sampling data from a stable visual world with accompanying context. This article presents a form of contrastive learning that uses spatial context from small image sets obtained from a simulated environment to train a model that generalizes to the classification of natural images and spatial perception, demonstrating improved performance compared to state-of-the-art contrastive methods. While this work is a modest step forward, the long-term ambition is to give intelligent agents a greater ability to learn on their own by exploring their surroundings. Intelligent agents with this capability have the potential to solve a diverse range of problems and make a significant impact on society.
keywords:
[…]
{highlights}
Created an approach that uses spatial context as a similarity signal
Developed a method for constructing image datasets using an environmental sampling agent
Training with contextual information improves state of the art in contrastive learning
Simulated data provide new forms of physically realistic augmentations
{graphicalabstract}
1 Summary
@@
2 Introduction
@@
Figure 1: The impact of position on the appearance, lighting, and camera distance/focal length of an image (A) The perspective of a room can greatly impact its appearance when rendered from different positions in the ThreeDWorld simulated environment. (B) The natural lighting of a scene can significantly alter its appearance when captured at different times of the day. Photos courtesy of Federico Adolfi. (C) The head and facial features of a statue may appear differently when captured with different focal lengths and camera distances. Photos courtesy of James Z. Wang.
@@
3 Results
3.1 Simulated datasets provide a source of spatial similarity
@@
Table 1: Comparison between the baseline and ESS on House environment with modified thresholds “Threshold” indicates that if a sample’s position and rotation difference relative to the key samples is below (x meters, y degrees), it is designated a positive sample. The column “Positive Pairs” shows the average number of positive samples in the dictionary for each threshold. N/A indicates that one or more thresholds were omitted from the similarity metric. “House14KLong” indicates that the number of pretext training epochs was increased to equate total training between the 14K and 100K datasets. $\uparrow$ denotes that higher values of this term are preferable. $\downarrow$ denotes that lower values of this term are more favorable. Numbers after $\pm$ represent the standard error of the mean, rounded to a minimum of .01. Numbers in bold highlight the best downstream classification results on each dataset.
Pretext Training Stage Positive Pretext Task Downstream ImageNet Classification
@@ 3.2 ESS has superiority over instance discrimination
3.3 Richer exploration of an environment improves learning
@@
3.4 ESS is complementary to other contrastive learning approaches
@@
3.5 Simulated lighting is a complementary augmentation
Table 5: Comparison of the baseline and ESS-MB trained on House100K on the room classification task for images from the House14K, House14KLighting and Apt14K datasets
@@ 3.5 Simulated lighting is a complementary augmentation
3.6 ESS training improves localization
In the spatial localization task, pretrained models are fine-tuned to estimate the position and rotation of the agent. As shown in Table 6, ESS-MB consistently achieves lower losses compared to the baseline for both datasets. Specifically, ESS-MB predicts the position of images with an error of under 1 meter and 2 meters for House14K and Apt14K, respectively. ESS-MB training leads to better predictive accuracy in position by 0.15 meters in House14K and 0.51 meters in Apt14K. While both models exhibit notable rotation errors, ESS-MB outperforms the baseline in both tasks, with a superiority of 16.26 and 7.05 degrees for House14K and Apt14K, respectively.
4 Discussion
These results provide clear evidence that incorporating spatial context in environmental sampling significantly improves the effectiveness of contrastive learning compared to methods using an equivalent number of training epochs on the same dataset. Both rotation and position are important for defining whether a pair of views is similar enough. Moreover, the magnitude of the threshold for spatial similarity influences the learning outcome; excessively large thresholds might mislabel highly distinct views as positive pairs. Additionally, we discovered that resampling the same environment to acquire more images substantially boosts downstream accuracy, even if the images originate from identical rooms with the same furnishings and largely similar trajectories. Collectively these findings support the ability of visual learning algorithms to efficiently extract visual pattern information from a given environment, both by tracking the history of spatial information and denser reexploration of the same locations from slightly different positions and view angles as exemplified in Figure 1A.
@@ 4 Discussion
Figure 3: Illustration of a more comprehensive approach to evaluating spatial similarity, which considers not only the distance and angle between two views but also the specific region of space being observed Even though the angular difference between the two views generated in (A) and (B), calculated as $\lvert\theta_{1}-\theta_{2}\rvert$, and the position difference are equivalent, the two views in (A) could be considered more similar due to their convergent perspective and shared focus on a specific region of space. In contrast, the views in (B) may be considered less similar due to their divergent perspective and lack of overlap in the region of space being observed.
@@
A limitation of our study is the modest overall accuracy achieved in the downstream image classification task. This is a predictable outcome given such a small set of images used in training and the narrow scope of the environment from which they were collected–especially when compared with the diversity of ImageNet. However, this limitation mimics the real-world learning scenarios experienced by embodied agents, such as children, who learn a robust basis set of visual representations through exposure to restricted environments. The evident gap between our current downstream accuracy and human performance in image classification suggests significant opportunities for improvement and future development in training algorithms that exploit environmental context. Given the relatively small size of the datasets used in our framework, there is potential for rapid experimentation and iterative refinement of similar algorithms. In light of these findings, we encourage the computer vision community to explore ways to narrow the accuracy gap for such datasets. We have made all of our datasets available on an explanatory website (see below).
5 Experimental Procedures
5.1 Resource availability
5.1.1 Lead contact
@@ 5.1.2 Materials availability
5.1.3 Data and code availability
Our dataset is based on the high-fidelity 3D virtual environment, ThreeDWorld,36 which can be downloaded at https://www.threedworld.org. Datasets used in this paper have been deposited at https://osf.io/w98gq/ under DOI 10.17605/OSF.IO/W98GQ and are publicly available as of the date of publication. All datasets are also available at http://www.child-view.com. We provide our two-stage dataset generation pipeline, along with the codes for conducting all the experiments and the pretraining and downstream checkpoints, at https://osf.io/ft59q/, under DOI 10.17605/OSF.IO/FT59Q and are publicly available as of the date of publication. Any additional information required to reanalyze the data reported in this paper is available upon request.
5.2 Dataset generation process
@@
One important advantage of varying light sources in a ray-traced virtual environment is its capacity to more accurately emulate the real-world physics of light reflection, resulting in a richer variety than basic augmentation techniques that merely shift spectral distributions. The ThreeDWorld platform features 95 distinct skyboxes as environment lighting conditions. We controlled an avatar to capture three images from the living room, stairs, and bedroom, maintaining consistent position and rotation in the House environment for each of the 95 skyboxes. The t-distributed stochastic neighbor embedding (t-SNE)48 was then used to cluster concatenations of those three images simulated under each of the 95 skyboxes. To explore lighting augmentations, we selected nine skyboxes, drawn from a $3\times 3$ grid of the t-SNE plot (Figure 4). A sample image from the House environment for each chosen skybox is shown within the t-SNE plot. Every image in the House14k and House100K datasets was generated ten times, one with the default lighting condition of ThreeDWorld and also one for each of these nine skyboxes. The resulting datasets are titled House14KLighting and House100KLighting. For more details on the lighting models within ThreeDWorld, readers can refer to the primary reference.36
@@
5.3 Implementation details of ESS-MB and ESS-MW
The difference between rotations $r_{i}$ and $r_{j}$ is defined as:
$\Delta_{\text{rot.}}=\min(|r_{i}-r_{j}|,360-|r_{i}-r_{j}|)\;.$ (2)
The loss function for image $i$ is then calculated as follows:
This strategy makes use of the spatial information from the environment to define the positive pairs. As there are often multiple samples in the dictionary that fall within the spatial similarity threshold relative to the query image, we use ESS-MB, where MB indicates there are multiple positive pairs. This strategy ensures that we don’t miss useful information or compromise training efficiency by only focusing on a single sample with high similarity to the query images. In ESS-MB, every sample within the spatial similarity threshold is treated as an equally valid positive pair during the calculation of loss, which is inspired by the supervised contrastive learning method.49 This approach allows for a more comprehensive consideration of relevant samples, leading to improved performance compared to using just a single positive pair.
@@ 5.4 Pretext training
5.5 Evaluation of the learned representations
The accuracy of the pretext task: The baseline model only considered one positive pair. The accuracy computation for the pretext training is different from ESS-MB which has multiple positive pairs and the two cannot be directly compared. In the baseline model, accuracy was calculated by determining if the pair with the highest cosine similarity was the pre-defined positive pair. For ESS-MB, accuracy was computed by applying the $sigmoid$ function to each cosine similarity score. If the result was greater than the threshold of 0.95, the pair was predicted to be positive; otherwise, it was predicted to be negative. The predicted result was then compared to the pre-defined positive pair according to the positions and rotation to calculate the accuracy.
@@ 5.5 Evaluation of the learned representations
ImageNet classification task: To evaluate the quality of the learned representations, as in MoCo, we added a linear classifier on top of the fixed backbone architecture and trained only the last added layer for 50 epochs of the ImageNet.
Room classification task: In this task, we trained a linear classifier to label a given image according to what room it had been generated in using the features from each pretrained model. Each downstream model was trained for 20 epochs. The House environment includes 8 rooms, while the Apt environment consists of 9 rooms. Each image is labeled with a number, ranging from 0 to 7 (for House) or up to 8 (for Apt), to represent the room where it was captured. The boundaries and illustrations of each room are included in Supplemental Information. In each dataset, $80\%$ of images were used for the training and the remaining $20\%$ for testing. In the House14KLighting dataset, Mosaic_tunnel and Venice_sunrise lighting conditions were only applied to the test data. Meanwhile, each training image was randomly assigned one of the other seven lighting conditions.
Spatial localization task: We added a single-layer neural network with four output nodes at the end of the pre-trained model. The training utilized $80\%$ of the images from each dataset, setting aside the remaining $20\%$ for testing purposes. In the spatial localization task, pretrained models are fine-tuned to estimate the position ($x_{\text{p}}$, $y_{\text{p}}$, $z_{\text{p}}$) and rotation $r_{p}$ of each image from the House14K and Apt14K datasets. The loss function, denoted as $L$, is defined by
@@
5.6 Hyperparameter sensitivity analysis
@@
Figure 7: Impact of varied pre-training hyperparameters on downstream test accuracy and pretext loss (A) The effect of hyperparameter variations on downstream test accuracy. (B) The effect of hyperparameter variations on pretext loss during pretraining. Regarding batch sizes, temperature parameters, thresholds, and dictionary sizes, blue indicates that the parameter value has been doubled, while red indicates that it has been halved. $*{}^{*}$ For lighting conditions, blue represents the default setting, whereas red represents the use of multiple skyboxes.
6 Acknowledgments
@@ 8 Declaration of Interests
9 Declaration of AI and AI-Assisted Technologies in the Writing Process
@@
Supplemental Information
Room boundaries in two environments
The Archviz House (‘House’) consists of eight distinct rooms, with example images depicted in Figure S1. These rooms are the kitchen, lower hall, lower bedroom, outer deck, upstairs piano room, bathroom, and upstairs bedroom. Images are classified based on their position inside the defined boundaries of each room. Certain images, such as those captured on the stairs, do not fall within the boundaries of any room and are therefore excluded from the evaluation. In House14K, 12,127 of 14,766 images are labeled. In House100K, 83,300 of 102,197 images are labeled. The boundaries and the number of samples for each room are listed in Table S1.
Figure S1: Eight rooms in the House environment (A) Kitchen. (B) Living room. (C) Lower hall. (D) Lower bedroom. (E) Outer deck. (F) Upstairs piano room. (G) Bathroom. (H) Upstairs bedroom.
Table S1: The boundaries of the eight rooms in the House environment and the number of samples for each room The coordinate ranges are measured in the ThreeDWorld virtual environment. “House14K images” and “House100K images” means the number of images in each category for House14K and House100K respectively.
Room name $x$-axis boundary $y$-axis boundary
Height
House14K
House100K
boundary
images
images
kitchen
(-17.00, -11.83)
(-1.48, 1.80)
(1.4, 3.5)
2079
7535
living room
(-17.00, -7.60)
(-7.00, -1.48)
(1.4, 5.2)
3225
32412
lower hall
(-6.30, -4.10)
(-4.30, -3.10)
(0.6, 3.5)
461
3490
lower bedroom
(-3.43, 0.05)
(-4.30, 1.60)
(0.6, 3.5)
1506
10155
outer deck
(-7.10, 6.00)
(-7.40, -4.75)
(0.4, 5.2)
2491
6721
upstairs piano room
(-3.20, -0.25)
(-4.30, -3.00)
(3.5, 5.2)
925
9882
bathroom
(-3.20, -0.25)
(-4.30, -3.00)
(3.5, 5.2)
419
2853
upstairs bedroom
( 0.00, 4.10)
(-4.30, 1.37)
(3.5, 5.2)
1021
10252
The Apartment (‘Apt’) layout consists of nine rooms, arranged in two rows. Rooms in the upper row in the floor plan are marked as rooms 0 to 4, from left to right. Rooms 5 to 8 are in the lower row in the same left-to-right sequence. The items placed in each room are carefully designed. For instance, entertainment facilities are filled in room 0. Rooms 3, 4, 6, and 7 serve as living rooms, each having a distinct style. Room 8 is a kitchen. In Apt14K, 9,855 of 14,487 images are labeled. Sample images and the specified boundaries for each room are shown in Figure S2 and Table S2, respectively.
Figure S2: Nine rooms in the Apt environment Letters A-I represent the rooms 0 through 8 in order.
Table S2: The boundaries of the nine rooms in the Apt environment and the number of samples for each room The coordinate ranges are measured in the ThreeDWorld virtual environment.
Room label
$x$-axis boundary
$y$-axis boundary
Images
0
(-10.6, -7.2)
( 1.45, 4.80)
680
1
(-7.2, -3.1)
( 1.45, 4.80)
659
2
(-3.1, 0.8)
( 1.45, 4.80)
1312
3
( 0.8, 6.8)
( 1.45, 4.80)
1295
4
( 6.8, 9.8)
( 1.45, 4.80)
749
5
(-10.6, -6,2)
(-5.70, -0.18)
1333
6
(-6.2, -3.1)
(-5.70, -0.18)
794
7
(-3.1, 3.3)
(-5.70, -0.18)
1269
8
( 3.3, 9.8)
(-5.70, -0.18)
1764
Learning rate comparison
In our pretext training, we adopted a learning rate of 0.3 instead of the suggested $0.015$ from MoCo V2. This adjustment was made based on its improved overall accuracy on our datasets. Table S3 shows the results of MoCo V2 with different learning rates when trained on the House100K dataset.
Table S3: Comparison results of MoCo V2 with two different learning rates trained on the House100K dataset
Learning Pretext Task Downstream ImageNet Classification
[…]
0.015 $3.73$ $73.45$ $209.25$ $7.61$
0.3 $4.43$ $82.11$ $4.71$ $17.36$
Details of the implementation of ESS-MB on other models
SimCLR31 is a popular contrastive learning model in which the positive pair of an augmented view is itself. Negative pairs are other augmented samples from the same batch. Our ESS-MB on SimCLR found positive samples from the same batch according to spatial information. All parameters were the same as those in SimCLR. We ran the experiment on a single GPU for 200 epochs as suggested by the code.
DCL38 removes the positive pairs’ effect on the denominator of InfoNCE loss. We implemented the updated loss function based on our original ESS-MB model for both DCL and ESS-MB with DCL. All the parameters were the same as the ESS-MB on MoCo.
CLSA39 categorizes augmentation operations into ‘strong’ and ‘weak’ and tries to align the feature distance distribution of views derived from these two augmentation types when finding the positive pairs from the weak augmented samples simultaneously. CLSA inherits the structure of MoCo. Based on the implementation of CLSA, our approach found positive pairs of an augmented view from the dictionary. In our experiment, we kept the hyperparameters of CLSA but modified the dictionary size and learning rate to match our original ESS-MB.
NNCLR40 computes similarity according to the proximity within a latent space generated by the encoder to contrastively learn representations from unlabeled images. The Lightly package was used to run NNCLR simulations with the ResNet-18 backbone. To ensure a fair comparison, we switched the backbone of ESS-MB on MoCo to Resnet-18 and trained both models for 200 epochs in the pretraining phase.
MoCo V342 applied the contrastive learning structure to the Vision Transformer43 backbone. The dictionary size is set to 4096 to align with the threshold of ESS-MB approach. We run both models for 200 epochs in the pretraining with 256 batch size.
Assessing the clustering of learned features
After training on the datasets, we applied t-SNE48 on the features for a subset of images from the corresponding datasets to determine whether the training produced clustering of features from spatially proximal images. We randomly selected approximately 10,000 images that were inside the room boundaries and took the features from the fully trained ResNet model as input to the t-SNE. In the resulting t-SNE space, each image’s features were labeled with a number (and a corresponding color), indicating the room of its origin.
The t-SNE visualizations generated for both the baseline and ESS-MB models trained on House and Apt environments are shown in Figures S3 and S4, respectively. Furthermore, we used the Silhouette Coefficient,50 Calinski-Harabasz index,51 and Davies-Bouldin index52 as metrics to assess the clustering quality of the t-SNE outputs. These results are shown in Table S4. Both results show a model trained on the more extensive House100K dataset exhibited a stronger capability in distinguishing features generated from different rooms compared to the one trained on the smaller House14K dataset. For all the models trained on three datasets, ESS-MB exhibited reduced clustering relative to the baseline model. This might be attributed to the ESS-MB training approach, which tends to group features of spatially proximal images together, regardless of room boundaries. In contrast, the baseline MoCo model relies only on instance discrimination. As a result, cluster boundaries for nearby locations in adjacent rooms would not be as distinct with ESS-MB.
Figure S3: The t-SNE results of the learned features in the House environment (A) ESS-MB on House14K. (B) ESS-MB on House100K. (C) Baseline on House14K. (D) Baseline on House100K.
Figure S4: The t-SNE results of the learned features in the Apt environment (A) ESS-MB on Apt14K. (B) Baseline on Apt14K.
Table S4: Evaluation of the learned features CH and DB stand for Calinski-Harabasz and Davies-Bouldin indices, respectively. An upward arrow indicates that a higher value for the respective index denotes more effective clustering, while a downward arrow implies the reverse.
Pretext dataset
Model
Silhouette\ $\uparrow$
CH index\ $\uparrow$
DB index\ $\downarrow$
House100K
Baseline
0.2394
4538.16
0.8552
House100K
ESS-MB
0.1437
2393.92
1.8488
House14K
Baseline
-0.0548
1187.78
5.5367
House14K
ESS-MB
0.0972
1549.44
9.8386
[…]

### v2

@@
Incorporating simulated spatial context information improves the effectiveness of contrastive learning models
Lizhen Zhu Data Science and Artificial Intelligence Area, College of Information Sciences and Technology, The Pennsylvania State University, University Park, PA, USA James Z. Wang Data Science and Artificial Intelligence Area, College of Information Sciences and Technology, The Pennsylvania State University, University Park, PA, USA Human-Computer Interaction Area, College of Information Sciences and Technology, The Pennsylvania State University, University Park, PA, USA Department of Communication and Media, School of Social Sciences and Humanities, Loughborough University, Loughborough, Leicestershire, UK Wonseuk Lee Department of Computer Science and Engineering, The Pennsylvania State University, University Park, PA, USA Brad Wyble Department of Psychology, The Pennsylvania State University, University Park, PA, USA
Summary
@@
Keywords
Contrastive learning, Virtual environment, Developmental psychology, Deep learning, Computer vision, Intelligent agent, Robotics, Childhood learning
Introduction
@@
Figure 1: The impact of position on the appearance, lighting, and camera distance/focal length of an image (A) The perspective of a room can greatly impact its appearance when rendered from different positions in the ThreeDWorld simulated environment. (B) The natural lighting of a scene can significantly alter its appearance when captured at different times of the day. Photos courtesy of Federico Adolfi. (C) The head and facial features of a statue may appear differently when captured with different focal lengths. Photos courtesy of James Z. Wang.
@@
Results
Simulated datasets provide a source of spatial similarity
@@ Results
ESS has superiority over instance discrimination
Table 1: Comparison between the baseline and ESS on House environment with modified thresholds “Threshold” indicates that if a sample’s position and rotation difference relative to the key samples is below (x meters, y degrees), it is designated a positive sample. The column “Pos. Pairs” shows the average number of positive samples in the dictionary for each threshold. N/A indicates that one or more thresholds were omitted from the similarity metric. “House14KLong” indicates that the number of pretext training epochs was increased to equate total training between the 14K and 100K datasets. $\uparrow$ denotes that higher values of this term are preferable. $\downarrow$ denotes that lower values of this term are more favorable. Numbers after $\pm$ represent the standard error of the mean, rounded to a minimum of .01. Numbers in bold highlight the best downstream classification results on each dataset.
Pretext Training Stage Pos. Pretext Task Downstream ImageNet Classification
@@
Richer exploration of an environment improves learning
@@
Simulated lighting is a complementary augmentation
@@
ESS training improves localization
Table 5: Comparison of the baseline and ESS-MB models trained on House100K on the room classification task for images from the House14K, House14KLighting and Apt14K datasets
@@
In the spatial localization task, pretrained models are fine-tuned to estimate the position and rotation of the agent. As shown in Table 6, ESS-MB consistently achieves lower losses compared to the baseline for both datasets. Specifically, ESS-MB predicts the position of images with an error of under 1 meter and 2 meters for House14K and Apt14K, respectively. ESS-MB training leads to better predictive accuracy in position by 0.15 meters in House14K and 0.51 meters in Apt14K.
@@
While both models exhibit notable rotation errors, ESS-MB outperforms the baseline in both tasks, with a superiority of 16.26 and 7.05 degrees for House14K and Apt14K, respectively.
Discussion
Interpretation of the results
Inspired by the processes of childhood learning, these results provide clear evidence that incorporating spatial context in environmental sampling significantly improves the effectiveness of contrastive learning compared to methods using an equivalent number of training epochs on the same dataset. Both rotation and position are important for defining whether a pair of views is similar enough. Moreover, the magnitude of the threshold for spatial similarity influences the learning outcome; excessively large thresholds might mislabel highly distinct views as positive pairs. Additionally, we discovered that resampling the same environment to acquire more images substantially boosts downstream accuracy, even if the images originate from identical rooms with the same furnishings and largely similar trajectories. Collectively these findings support the ability of visual learning algorithms to efficiently extract visual pattern information from a given environment, both by tracking the history of spatial information and denser reexploration of the same locations from slightly different positions and view angles as exemplified in Figure 1A.
@@ Discussion
Limitations of the study
A limitation of our study is the modest overall accuracy achieved in the downstream image classification task. This is a predictable outcome given such a small set of images used in training and the narrow scope of the environment from which they were collected–especially when compared with the diversity of ImageNet. However, this limitation mimics the real-world learning scenarios experienced by embodied agents, such as children, who learn a robust basis set of visual representations through exposure to restricted environments. The evident gap between our current downstream accuracy and human performance in image classification suggests significant opportunities for improvement and future development in training algorithms that exploit environmental context. Given the relatively small size of the datasets used in our framework,
Figure 3: Illustration of a more comprehensive approach to evaluating spatial similarity, which considers not only the distance and angle between two views but also the specific region of space being observed Even though the angular difference between the two views generated in (A) and (B), calculated as $\lvert\theta_{1}-\theta_{2}\lvert$, and the position difference are equivalent, the two views in (A) could be considered more similar due to their convergent perspective and shared focus on a specific region of space. In contrast, the views in (B) may be considered less similar due to their divergent perspective and lack of overlap in the region of space being observed.
there is potential for rapid experimentation and iterative refinement of similar algorithms. In light of these findings, we encourage the computer vision community to explore ways to narrow the accuracy gap for such datasets. We have made all of our datasets available on an explanatory website (see Experimental procedures).
Future directions
@@
Experimental procedures
Resource availability
Lead contact
@@
Data and code availability
Our dataset is based on the high-fidelity 3D virtual environment, ThreeDWorld,36 which can be downloaded at https://www.threedworld.org. Datasets used in this paper have been deposited at the OSF under DOI 10.17605/OSF.IO/W98GQ and are publicly available as of the date of publication.48 All datasets are also available at http://www.child-view.com. We provide our two-stage dataset generation pipeline, along with the codes for conducting all the experiments and the pretraining and downstream checkpoints, at the OSF, under DOI 10.17605/OSF.IO/FT59Q and are publicly available as of the date of publication.49 Any additional information required to reanalyze the data reported in this paper is available upon request.
Dataset generation process
@@
One important advantage of varying light sources in a ray-traced virtual environment is its capacity to more accurately emulate the real-world physics of light reflection, resulting in a richer variety than basic augmentation techniques that merely shift spectral distributions. The ThreeDWorld platform features 95 distinct skyboxes as environment lighting conditions. We controlled an avatar to capture three images from the living room, stairs, and bedroom, maintaining consistent position and rotation in the House environment for each of the 95 skyboxes. The t-distributed stochastic neighbor embedding (t-SNE)50 was then used to cluster concatenations of those three images simulated under each of the 95 skyboxes. To explore lighting augmentations, we selected nine skyboxes, drawn from a $3\times 3$ grid of the t-SNE plot (Figure 4). A sample image from the House environment for each chosen skybox is shown within the t-SNE plot. Every image in the House14k and House100K datasets was generated ten times, one with the default lighting condition of ThreeDWorld and also one for each of these nine skyboxes. The resulting datasets are titled House14KLighting and House100KLighting. For more details on the lighting models within ThreeDWorld, readers can refer to the primary reference.36
@@
Implementation details of ESS-MB and ESS-MW
The difference between rotations $r_{i}$ and $r_{j}$ is defined as:
$\Delta_{\text{rot.}}=\min(|r_{i}-r_{j}|,360-|r_{i}-r_{j}|)\;.$ (2)
@@
The loss function for image $i$ is then calculated as follows:
This strategy makes use of the spatial information from the environment to define the positive pairs. As there are often multiple samples in the dictionary that fall within the spatial similarity threshold relative to the query image, we use ESS-MB, where MB indicates there are multiple positive pairs. This strategy ensures that we don’t miss useful information or compromise training efficiency by only focusing on a single sample with high similarity to the query images. In ESS-MB, every sample within the spatial similarity threshold is treated as an equally valid positive pair during the calculation of loss, which is inspired by the supervised contrastive learning method.51 This approach allows for a more comprehensive consideration of relevant samples, leading to improved performance compared to using just a single positive pair.
@@
Evaluation of the learned representations
The accuracy of the pretext task
The baseline model only considered one positive pair. The accuracy computation for the pretext training is different from ESS-MB which has multiple positive pairs and the two cannot be directly compared. In the baseline model, accuracy was calculated by determining if the pair with the highest cosine similarity was the pre-defined positive pair. For ESS-MB, accuracy was computed by applying the $sigmoid$ function to each cosine similarity score. If the result was greater than the threshold of 0.95, the pair was predicted to be positive; otherwise, it was predicted to be negative. The predicted result was then compared to the pre-defined positive pair according to the positions and rotation to calculate the accuracy. The pretraining accuracies of the main experiments from Table 1 are shown in Table 8.
ImageNet classification task
@@
To evaluate the quality of the learned representations, as in MoCo, we added a linear classifier on top of the fixed backbone architecture and trained only the last added layer for 50 epochs of the ImageNet.
Room classification task
In this task, we trained a linear classifier to label a given image according to what room it had been generated in using the features from each pretrained model. Each downstream model was trained for 20 epochs. The House environment includes 8 rooms, while the Apt environment consists of 9 rooms. Each image is labeled with a number, ranging from 0 to 7 (for House) or up to 8 (for Apt), to represent the room where it was captured. The boundaries and illustrations of each room are included in Supplemental Information. In each dataset, $80\%$ of images were used for the training and the remaining $20\%$ for testing. In the House14KLighting dataset, Mosaic_tunnel and Venice_sunrise lighting conditions were only applied to the test data. Meanwhile, each training image was randomly assigned one of the other seven lighting conditions.
Spatial localization task
We added a single-layer neural network with four output nodes at the end of the pre-trained model. The training utilized $80\%$ of the images from each dataset, setting aside the remaining $20\%$ for testing purposes. In the spatial localization task, pretrained models are fine-tuned to estimate the position ($x_{\text{p}}$, $y_{\text{p}}$, $z_{\text{p}}$) and rotation $r_{p}$ of each image from the House14K and Apt14K datasets. The loss function, denoted as $L$, is defined by
@@
Hyperparameter sensitivity analysis
@@
Figure 7: Impact of varied pre-training hyperparameters on downstream test accuracy and pretext loss (A) The effect of hyperparameter variations on downstream test accuracy. (B) The effect of hyperparameter variations on pretext loss during pretraining. Regarding batch sizes, temperature parameters, thresholds, and dictionary sizes, blue indicates that the parameter value has been doubled, while red indicates that it has been halved. For lighting conditions, blue represents the default setting, whereas red represents the use of multiple skyboxes.
@@
Acknowledgments
@@
Declaration of AI and AI-Assisted Technologies in the Writing Process